which of the following bonds is the most polar? a.H-C1 b.C-C1 c.H-H d.N-F


1
Expert's answer
2022-03-28T17:07:07-0400

Which bond has maximum value of electronegativity difference that bond has most polar.

So, H-Cl bond has the maximum value of electronegativity difference, and the value is 0.96, so HCl bond has most polar bond on compare to other bond.

Electronegativity value of H, Cl, C, N, F

H=2.20, Cl=3.16, C=2.55, N=3.04, F=3.98

Calculate the maximum electronegativity difference of the following bond

(a) H-Cl = 3.16-2.20=0.96

(b) C-Cl = 3.16-2.55=0.61

(C) H-H = 2.20-2.20=0

(d) N-F = 3.98-3.04=0.94

So, on the basis of above calculation we can get the H-Cl bond has maximum value of electronegativity difference and the value is 0.96, so

H-Cl ,bond has most polar bond.